In hou many ways 5 sportswomen can be selected from a group of 10?
Answer by jim_thompson5910(35256)

C(n,r) = (n!)/(r!(n-r)!)

C(10,5) = (10!)/(5!*(10-5)!)

C(10,5) = (10!)/(5!*5!)

C(10,5) = (10*9*8*7*6*5!)/(5!*5!)

C(10,5) = (10*9*8*7*6)/(5!)

C(10,5) = (10*9*8*7*6)/(5*4*3*2*1)

C(10,5) = (30240)/(120)

C(10,5) = 252

So there are 252 ways. This is assuming that order doesn't matter.